zayy:

A small fruit basket with 6 apples and 6 oranges costs $7.50. A different fruit basket with 10 apples and 5 oranges costs $8.75. If x is the cost of one apple and y is the cost of one orange, the system of equations below can be used to determine the cost of one apple and one orange. 6x+6y=7.50 10x+5y=8.75

XTHb:

so there is this system of equations 6x +6y = 7,50 10x +5y = 8,75 so now just you need solve this system of equations - depend what method you wan need using substitution or elimination ?

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @XTHb so there is this system of equations 6x +6y = 7,50 10x +5y = 8,75 so now just you need solve this system of equations - depend what method you wan need using substitution or elimination ? \(\color{#0cbb34}{\text{End of Quote}}\) good idea - congrats !